OpenStudy (anonymous):

Solve -6=3x-6y 4x=4+5y for x and y or no solution and infinite

OpenStudy (amistre64):

3 -6 | -6 1 -2 | -2 (4) -4 5 |-4 4 -8 | -8 -4 5 |-4 --------- -3|-12 1|4 ; y = 4

OpenStudy (anonymous):

6,4

OpenStudy (amistre64):

1 -2 | -2 0 1 | 4 (2) 1 -2 | -2 0 2 | 8 --------- 1 0 | 6 ; x = 6

OpenStudy (anonymous):

oh look it is (6,4) again!

OpenStudy (anonymous):

4x -5y= 4 -3x+6y = 6 12 x- 15y = 12 -12x +24y = 24 9y = 36 y = 4 x = 4 +20 / 4 x = 24 /4 = 6

OpenStudy (amistre64):

6 and 4 are transcendant numbers lol

OpenStudy (anonymous):

u r too fast for me amis!!

OpenStudy (amistre64):

;)

OpenStudy (amistre64):

'sok; looks like gian is pulling up behind

OpenStudy (anonymous):

so every1 in an unseen race with amistre;)

OpenStudy (anonymous):

i was going to quit when i got to 1000. now i am going to quit when amistre gets there

OpenStudy (anonymous):

:D

OpenStudy (anonymous):

yeah looks like lol

OpenStudy (anonymous):

gian isn't giving up!!

OpenStudy (anonymous):

-6=3x-6y this can be written 6y = 3x + 6 (divide by 6) y = 1/2x +1 4 x=4+5y (substitute above value of y into this equation) 4x = 4 + 5(1/2x +1) expand brackets now 4x = 4 + 5/2x + 5 (subtract 5/2x from both sides and simplify RHS 3/2x = 9 multiply by 2 on both sides 3x = 18 now divide by 3 on both sides x = 6 Now sub this value of x into equation 1 -6 = 3(6) - 6y -6 = 18 - 6y now -6 -18 = -6y -24 = -6y y = 4 Phew we got there!

OpenStudy (anonymous):

well done

OpenStudy (anonymous):

Please no congratulations it makes me blush!